In order join a tennis class , you pay a $200 annual fee, then $10 for each class you go to . What is the average cost per class if you go to 10 classes
1
Expert's answer
2020-11-26T15:55:40-0500

Total cost = 200 + extra cost paid for the 10 classes

= 200 + 10 + 10 + 10 + 10 +10 + 10 +10 + 10 + 10 + 10 = $300

Average cost of the 10 classes = Total cost / number of classes

= 300/10 = $30

Therefore, the average cost if you go to 10 classes = $30.